carlosbigbain:

how are 50 and 1 over 2 Division sign 1 over 4 equals 202 all related

carlosbigbain:

plz someone answer i really need this answer

4evablood:

"""""In Multiplication, the 50 1/2 and 4 are factors, Four students each at fifty and one half tables. 202 is the product"""""""

carlosbigbain:

thank you plz more people answer cause i want the best answer i can get

jhonyy9:

\[(50 \frac{ 1 }{ 2 } ) / (1/4) = 202\] 50 *(1/2) = 50/2 = 25 25 / (1/4) = 25 *4 = 100 - i think you missed details from there

carlosbigbain:

"Give a verbal explanation that describes how these three numbers are related" thats what it says

carlosbigbain:

i kinda still need help